OpenStudy (anonymous):

find the slope-intercept equation of the line that has the given characteristics slope 0 and y intercept (0,3)

OpenStudy (ash2326):

Slope Intercept form of a line \[y=mx+c\] m= slope of the line c= y intercept Just plugin the numbers and you'll get the line's equation

OpenStudy (anonymous):

so the equation should read y=0x+(0,3)

OpenStudy (ash2326):

nope, just y part of the intercept \[(0, 3)=> 3\] \[y=0\times x +3\]

OpenStudy (anonymous):

so equation should read y=0x+3

OpenStudy (ash2326):

yeah, do you understand?

OpenStudy (anonymous):

yea now i do

OpenStudy (anonymous):

thanks

OpenStudy (ash2326):

:D

OpenStudy (anonymous):

so an equation for m=3/4,(3,-2) is y=3/4x+-2

OpenStudy (ash2326):

nope, if x=0 then only it's a y-intercept, to find y intercept or c \[y=mx+c\] We have m with us \[y=\frac 34 x+c\] to find c, put the given point (3, -2) \[-2=\frac 34 \times 3+c\] Now find c and then write the equation!!

OpenStudy (anonymous):

the ? reads find an equation of the line having the given slope and containing the given point m=3/4,(3,-2)

OpenStudy (anonymous):

and the answer starts out with y=

OpenStudy (ash2326):

find c and then plugin the numbers in \[y=mx+c\]

OpenStudy (anonymous):

i dont understand what you mean

OpenStudy (anonymous):

cause so far all i have is y=3/4x+b

OpenStudy (ash2326):

The equation is give as \[y=mx+c\] We have m given as 3/4 and we need to find c \[y=\frac 3 4x+c...................eq(1)\] for that use the point (3, -2) Since the line passes through (3, -2). This point will satisfy the line's equation put x=3 and y=-2 in eq (1) and find c

OpenStudy (anonymous):

so next step is -2=2.25+c

OpenStudy (anonymous):

and then -2.25-2=-2.25+2.25+c

OpenStudy (anonymous):

then answer should 4.25=c